### v3

# Gradient-based Sample Selection for Faster Bayesian Optimization

Abstract
Bayesian optimization (BO) is an effective technique for black-box optimization. However, its applicability is typically limited to moderate-budget problems due to the cubic complexity of fitting the Gaussian process (GP) surrogate model. In large-budget scenarios, directly employing the standard GP model faces significant challenges in computational time and resource requirements. In this paper, we propose a novel approach, gradient-based sample selection Bayesian Optimization (GSSBO), to enhance the computational efficiency of BO. The GP model is constructed on a selected set of samples instead of the whole dataset. These samples are selected by leveraging gradient information to remove redundancy while preserving diversity and representativeness. We provide a theoretical analysis of the gradient-based sample selection strategy and obtain explicit sublinear regret bounds for our proposed framework. Extensive experiments on synthetic and real-world tasks demonstrate that our approach significantly reduces the computational cost of GP fitting in BO while maintaining optimization performance comparable to baseline methods.
Keywords Bayesian optimization, Large-scale, Gaussian process, Gradient information, Subset selection, Sublinear regret bound, Faster
1 Introduction
Bayesian optimization (BO) [Frazier, 2018] is a successful approach to black-box optimization that has been applied in a wide range of applications, such as hyperparameter optimization and mineral resource exploration. BO’s strength lies in its ability to represent the unknown objective function through a surrogate model and by optimizing an acquisition function [Garnett, 2023, Wang et al., 2023]. BO consists of a surrogate model, which provides a global prediction for the unknown objective function, and an acquisition function that serves as a criterion to determine the next sample to evaluate. In particular, the Gaussian process (GP) model is often preferred as the surrogate model due to its versatility and reliable uncertainty estimation. However, the GP model often suffers from large data sets, making it more suitable for small-budget scenarios [Binois and Wycoff, 2022]. To fit a GP model, the dominant complexity in computing the inversion of the covariance matrix is $\mathcal{O}(n^{3})$, where $n$ is the number of data samples. As the sample set grows, the computational burden increases substantially. This limitation poses a significant challenge for scaling BO to real-world problems with large sample sets.
Figure 1: Illustration of GP fitting with sample selection. Left: GP fitted with 10 samples. Right: GP fitted with 6 selected samples. With fewer selected samples, we can still fit a good GP to estimate the black box function, guiding us in finding the global optimum.
Despite the various approaches to improve the computational efficiency of BO, including parallel BO [González et al., 2016, Daulton et al., 2021, 2020, Eriksson et al., 2019], kernel approximation [Kim et al., 2021, Jimenez and Katzfuss, 2023, Hensman et al., 2013, Williams and Seeger, 2000] and sparse GP [Lawrence et al., 2002, Leibfried et al., 2020, McIntire et al., 2016], the computational overhead remains a burden in practice [Shahriari et al., 2015]. Kernel approximation methods typically involve simplifying or approximating the kernel matrix, which can degrade accuracy and lead to suboptimal performance if the approximation is too coarse. Sparse GP methods, while reducing computational complexity, introduce additional complexity through corrective terms to maintain approximation accuracy, and existing iterative implementations in Sparse GP [McIntire et al., 2016] frequently involve adding new samples while removing previous ones, potentially causing inefficiencies or suboptimal sample selection in complex optimization landscapes.
During the iterative search process of BO, some samples can become redundant and contribute little to the additional information gain. Such samples collected in earlier stages thus diminish in importance as the process evolves. For instance, excessive searching around identified minima becomes redundant once the optimal value has been determined, as these samples cease to offer meaningful insights for further optimization. To efficiently fit a GP, it is essential to focus on samples that provide the most informative contributions. As shown in Figure 1, carefully selected samples can effectively fit a GP. Despite the reduced number of samples, the GP still captures the key trends and features of the true function while maintaining reasonable uncertainty bounds. In this paper, we propose incorporating the gradient-based sample selection technique into the BO framework to enhance its scalability and effectiveness in large-budget scenarios. This technique originated in continual learning with online data stream [Aljundi et al., 2019]. The previous data are selectively sampled and stored in a replay buffer to prevent catastrophic forgetting and enhance model fitting. By using gradient information to gauge the value of each sample, one can more judiciously decide which samples are most essential for building a subset, and maintain the most representative subset of BO samples. This subset is then used to fit the GP model, accelerating the BO process while ensuring efficient and effective GP fitting. To the best of our knowledge, this work is the first to propose using gradient information for subset selection to accelerate Bayesian optimization. We summarize our main contributions as follows:
  • Efficient computations. We propose Gradient-based Sample Selection Bayesian Optimization (GSSBO) that addresses the scalability challenges associated with large-budget scenarios. Our approach is an out-of-the-box algorithm that can seamlessly integrate into existing BO frameworks with only a small additional computational overhead.
  • Theoretical analysis. We provide a rigorous theoretical analysis of the regret bound for the GSSBO. Theoretical results show that the regret bound of our proposed algorithm is similar to that of the standard GP-UCB.
  • Empirical validations. We conduct comprehensive numerical experiments, including synthetic and real-world test problems, to demonstrate that compared to baseline methods. The proposed algorithm achieves comparable performance and significantly reduces computational costs. These results verify the benefit of using gradient information to select a representative subset of samples.
2 Related Works
BO with Resource Challenges. In practical applications, BO faces numerous challenges, including high evaluation costs, input-switching costs, resource constraints, and high-dimensional search spaces. Researchers have proposed a variety of methods to address these issues. For instance, parallel BO employs batch sampling to improve efficiency in large-scale or highly concurrent scenarios [González et al., 2016, Daulton et al., 2021, 2020, Eriksson et al., 2019]. Kernel approximation methods, such as random Fourier features, map kernels onto lower-dimensional feature spaces, thus accelerating kernel-based approaches [Rahimi and Recht, 2007, Kim et al., 2021]. Multi-fidelity BO leverages coarse simulations with a limited number of high-fidelity evaluations to reduce the overall cost [Kandasamy et al., 2016]. For high-dimensional tasks, random embeddings or active subspaces help reduce the search dimensionality [Wang et al., 2016, Nayebi et al., 2019]. Meanwhile, sparse GP significantly reduces computational complexity by introducing inducing points [Lawrence et al., 2002, Leibfried et al., 2020, McIntire et al., 2016, Moss et al., 2023]. However, these approaches face limitations in practical scenarios and usually sacrifice performance for scaling. Calandriello et al. [Calandriello et al., 2022] scaling GP optimization by repeatedly evaluating each selected point until its posterior uncertainty falls below a preset threshold, thus limiting the number of datasets. However, the dataset still grows with time, and the algorithm’s dependence on its initial sample set means that low-value points selected early on remain permanently in the model, potentially inflating computational overhead.
BO with Gradient Information. The availability of derivative information can significantly simplify optimization problems. Ahmed et al. [Ahmed et al., 2016] highlights the potential of incorporating gradient information into BO methods and advocates for its integration into optimization frameworks. Wu et al. [Wu and Frazier, 2016] introduced the parallel knowledge gradient method for batch BO, achieving faster convergence to global optima. Then they introduce d-KG [Wu et al., 2017] as a new acquisition function, which systematically introduces gradient information into the BO and proves that using gradient information can strictly improve the information value. Rana et al. [Rana et al., 2017] incorporated GP priors to enable gradient-based local optimization. Chen et al. [Chen et al., 2018] proposed a unified particle-optimization framework using Wasserstein gradient flows for scalable Bayesian sampling. Bilal et al. [Bilal et al., 2020] demonstrated that BO with gradient-boosted regression trees performed well in cloud configuration tasks. Tamiya et al. [Tamiya and Yamasaki, 2022] developed stochastic gradient line BO (SGLBO) for noise-robust quantum circuit optimization. Penubothula et al. [Penubothula et al., 2021] funded local critical points by querying where the predicted gradient is zero. Zhang et al. [Zhang and Rodgers, 2024] introduced BO of gradient trajectory (BOGAT) for efficient imaging optimization. Makrygiorgos et al. [Makrygiorgos et al., 2025] integrate exact gradient observations into the Bayesian neural network surrogate’s training loss. Although these methods leverage gradient information to improve optimization efficiency and performance, they mainly focus on refining the GP model or acquisition functions.
Subset Selection. Subset selection is a key task in fields such as regression, classification, and model selection, aiming to improve efficiency by selecting a subset of features or data. Random subset selection, a simple and widely used method, involves randomly sampling data, often for cross-validation or bootstrap [Hastie, 2009]. Importance-based selection focuses on high-value data points, while active learning targets samples that are expected to provide the most information, improving model learning [Quinlan, 1986]. Filter methods rank features using statistical measures such as correlation or variance, selecting the top-ranked ones for modeling [Guyon and Elisseeff, 2003]. Narendra et al. [Narendra and Fukunaga, 1977] introduced a branch-and-bound algorithm for efficient feature selection. Yang et al. [Yang et al., 2022] proposed dataset pruning, an optimization-based sample selection method that identifies the smallest subset of training data to reduce training costs. Ash et al. [Ash et al., 2019] employs the k-means++ algorithm in the gradient space for diversity sampling in active learning. Oglic et al. [Oglic and Gärtner, 2017] first maps each data point into the reproducing kernel Hilbert space (RKHS), then uses a max–min coverage strategy in the RKHS to sequentially sample $K$ landmarks, which are employed to construct the Nyström low-rank approximation. This method is dependent on the quality of the selected landmarks. Hayakawa et al. [Hayakawa et al., 2023] provides tighter expected error bounds under a continuous measure for the same underlying idea. However, computing the Mercer decomposition in high dimensions incurs substantial computational cost and suffers from severe error degradation. Because both of these methods perform their approximations in RKHS at high computational expense, they are most suitable for offline, batch-mode resampling. Zhu et al. [Zhu, 2016] proposed a “pilot estimate” to approximate the gradient of the objective function. The core idea is to compute the gradient information corresponding to each data point based on an initial parameter estimate and identify data points with larger gradient values as more “important” samples for subsequent optimization. Despite these advancements, directly applying subset selection methods to BO often yields suboptimal results, necessitating further exploration to integrate subsampling effectively into BO frameworks.
3 Preliminaries
3.1 Bayesian Optimization and Gaussian Processes
BO aims to find the global optimum $x^{*}\in\mathcal{X}$ of an unknown reward function $f:\mathcal{X}\to\mathbb{R}$, over the $n$-dimensional input space $\mathcal{X}=[0,1]^{n}$. Throughout this paper, we consider maximization problems, i.e., we aim to find $x^{*}\in\mathcal{X}$ such that $f(x^{*})\geq f(x)$ for all $x\in\mathcal{X}$, get the optimal point $x^{*}=\arg\max_{x\in\mathcal{X}}f(x)$ as quickly as possible. GPs are one of the fundamental components in BO, providing a theoretical framework for modeling and prediction in a black-box function. In each round, a sample $x_{t}$ is selected based on the current GP’s posterior and acquisition function. The observed values $y_{t}$ and $x_{t}$ are then stored in the sample buffer, and the GP surrogate is updated according to these samples. This iterative process of sampling and updating continues until the optimization objectives are achieved or the available budget is exhausted. The key advantage of GPs lies in their nonparametric nature, allowing them to model complex functions without assuming a specific form. GPs are widely used for regression (Gaussian Process Regression [Schulz et al., 2018], GPR) and classification tasks due to their flexibility and ability to provide uncertainty estimates. Formally, a GP can be defined as: $f(\mathbf{x})\sim\mathcal{GP}(\mu(x),k(x,x^{\prime}))$, where $\mu(x)$ is the mean function, often assumed to be zero, and $k(x,x^{\prime})$ is the covariance function, defining the similarity between points $x$ and $x^{\prime}$. It should be noted that the algorithmic complexity of GP updates is $\mathcal{O}(n^{3})$, where $n$ is the number of observed samples. As the sample set grows, the computational resources required for these updates can become prohibitively expensive, especially in large-scale optimization problems.
3.2 Diversity-based Subset Selection
Due to limited computing resources, properly selecting samples instead of using all samples to fit a model is more efficient in problems with a large sample set. In continual learning, this helps overcome the catastrophic forgetting of previously seen data when faced with online data streams. Suppose that we have a model fitted on observed samples $\mathcal{D}\triangleq\{(x_{1},y_{1}),\ldots,(x_{t},y_{t})\}$, where $x_{i}\in\mathcal{X}$ and $y_{i}$ is the corresponding observation. In the context of subset selection, our objective is to ensure that each newly added sample contributes meaningfully to the optimization process. That is, a constraint ensures that when we select new samples for the sample subset, the performance of the model after the new samples are added will not be worse than the performance of the previous subset samples. Let $\mathbf{g}_{t}$ be the gradient of the sample at time $t$. Following [Aljundi et al., 2019], we rephrase the constraints with respect to the gradients. Specifically, the constraint can be rewritten as $\langle\mathbf{g}_{t},\mathbf{g}_{i}\rangle\geq 0,\forall i\in\{1,\dots,n-1\}.$ This transformation simplifies the constraint by focusing on the inner product of the gradients, which are nonnegative, such that there will be no performance degradation. To solve the constraint, we can use the geometric properties of the gradients. Note that optimizing the solid angle subtended by the gradients is computationally expensive. According to the derivation in [Aljundi et al., 2019], the sample selection problem is equivalent to maximizing the variance of the gradient direction of the samples in the fixed-size buffer. By maximizing the variance of the gradient directions, we ensure that the selected samples represent diverse regions, and therefore the buffer contains diverse samples, each contributing unique information to the optimization process. How to determine the buffer size will be detailed in Section 4.3. The previous problem thus becomes a surrogate for selecting a subset $\mathcal{U}$ of the samples that maximizes the diversity of their gradients:
$\text{Var}_{\mathcal{U}}\left[\frac{\mathbf{g}}{\|\mathbf{g}\|}\right]=1-\frac{1}{M^{2}}\sum_{i,j\in\mathcal{U}}\frac{\langle\mathbf{g}_{i},\mathbf{g}_{j}\rangle}{\|\mathbf{g}_{i}\|\|\mathbf{g}_{j}\|}.$ (1)
Here, $M$ denotes the buffer size and $\textbf{g}/\|\textbf{g}\|$ is the normalized gradient vector. The larger the value of this formula, the more dispersed the gradient direction of the selected samples is (the higher the diversity is). The reformulated problem (1) transforms the sample selection process from a sequential approach (adding samples to the subset one at a time) into a batch selection approach (samples are selected all at once). This empirical surrogate objective is agnostic to how gradient information is computed, making it straightforward to integrate into subset‐based methods.
4 Bayesian Optimization with Gradient-based Sample Selection
4.1 Gradient Information from GP
In the previous section, we introduced how to use the gradient to select a diverse subset of samples. However, within the Bayesian optimization framework using Gaussian processes, we do not have access to second-order information with respect to the input $x_{i}$, nor an explicit differentiable form of the objective to build $x_{i}$-gradients or sensitivities. We are more concerned with the impact on the GP posterior of the information carried by the observation $y_{i}$. Therefore, in light of the pilot estimate-based gradient information acquisition method in [Zhu, 2016], we propose a new method for gradient information acquisition in GPs: $g_{i}\;=\;\frac{\partial}{\partial y_{i}}\,\log p(\mathbf{y}\mid\mathbf{X},\theta).$ We focus on the gradient of output $\mathbf{y}$ rather than input $\mathbf{X}$ as the derivative of $\mathbf{y}$ is simpler and directly reflects each sample’s contribution. In a GP model, given a set of samples $\mathcal{D}=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ that follows a multivariate normal distribution with mean $\mu$ and covariance matrix $\mathbf{K}$, where $\mathbf{K}$ is constructed from a kernel function $k(\mathbf{x}_{i},\mathbf{x}_{j};\theta)$ and $\theta$} represents the hyperparameters, the probability density function of a multivariate Gaussian distribution is $p(\mathbf{y}|\mathbf{X},{\theta})=\frac{1}{(2\pi)^{n/2}|\mathbf{K}|^{1/2}}\exp\left(-\frac{1}{2}\mathbf{(y-\mu)}^{T}\mathbf{K}^{-1}\mathbf{(y-\mu)}\right)$. Taking the logarithm of it, we derive the log-likelihood function:
$\begin{split}\log p(\mathbf{y}|\mathbf{X},{\theta})=&-\frac{1}{2}\mathbf{(y-\mu)}^{T}\mathbf{K}^{-1}\mathbf{(y-\mu)}\\ &-\frac{1}{2}\log|\mathbf{K}|-\frac{n}{2}\log(2\pi).\end{split}$ (2)
Remark 4.1.
The log-likelihood function in (2) comprises three terms. The first term, $\frac{1}{2}\mathbf{(y-\mu)}^{T}\mathbf{K}^{-1}\mathbf{(y-\mu)}$, represents the sample fit under the covariance structure specified by $\mathbf{K}$. The second term, $\frac{1}{2}\log|\mathbf{K}|$, penalizes model complexity through the log-determinant of the covariance matrix. The third term, $\frac{n}{2}\log(2\pi)$, is a constant to the parameters and thus does not affect the gradient calculation.
The derivative of log likelihood with respect to $\mathbf{y}$ directly measures how sensitive this log-likelihood is to each observation $y_{i}$. Intuitively, if changing $y_{i}$ significantly alters the value of (2), that sample has a large marginal contribution to the fit. Hence, in subset selection schemes, one can use these gradient magnitudes to gauge how important each sample is, potentially adjusting their weights or deciding which samples to retain in a subset. To define the gradient for each sample, we first quantify each sample’s contribution to the log-likelihood. Since the second and third terms in (2), $\frac{1}{2}\log|\mathbf{K}|$ and $\frac{n}{2}\log(2\pi)$, do not depend on $\mathbf{y}$, both have no contribution to the gradient. Consequently, the gradient of the log-likelihood with respect to $\mathbf{y}$ is given by: $\frac{\partial\log p(\mathbf{y|X},{\theta})}{\partial\mathbf{y}}=-\mathbf{K}^{-1}(\mathbf{y}-\mu)$. Note that the $i$-th component, $\big(\mathbf{K}^{-1}(\mathbf{y}-\mu)\big)_{i}$, corresponds to the partial derivative of the log-likelihood with respect to $\mathbf{y}_{i}$; we denote this scalar sensitivity by $g_{i}$. Thus, we further define the vector embedding for each sample $i$,
$\mathbf{g}_{i}\;\triangleq\;\frac{\partial}{\partial y_{i}}\!\left(\frac{\partial\log p(\mathbf{y|X},\theta)}{\partial\mathbf{y}}\right)=-\,\mathbf{K}^{-1}\mathbf{e}_{i},$ (3)
where $\mathbf{e}_{i}$ is the $i$-th standard basis vector. After obtaining the gradient information of each sample, we can use the maximization of the variance of the angle between the two gradients as an alternative goal (i.e. maintaining the diversity of gradient directions). This gradient calculation is computationally efficient as the value of $\mathbf{K}^{-1}$ is available while updating the GP. Furthermore, the complexity of the additional computational burden introduced by the gradient calculation is $O(n^{2})$, and it is negligible compared to the complexity of GP updates (which is $O(n^{3})$), especially when $n$ is large.
4.2 Gradient-based Sample Selection
As the number of observed samples increases, fitting a GP model can become prohibitively expensive, especially in large-scale scenarios. A common remedy is to work with a subset of samples of size $M\ll N$, thereby reducing the computational cost of GP updates. The efficiency and effectiveness of GP model fitting in BO are closely related to the quality of the chosen subset. This raises the question: How do we choose a subset that remains representative and informative? Inspired by the success of gradient-based subset selection methods in machine learning, we propose leveraging gradient information to guide the selection of such subsets within BO. To this end, we introduce a gradient-based sample selection methodology to ensure representativeness within a limited sample buffer size. By harnessing gradient information, our approach maintains a carefully chosen subset of samples that not only eases computational burdens, but also preserves model quality, even as the sample set size grows. We begin by modeling the objective function $f$ with a GP and setting a buffer size $M$. Initially, the algorithm observes $f$ at $n_{0}$ samples, retaining these initial samples to preserve global information critical to the model. After each subsequent evaluation, if the number of samples exceeds $M$, we perform a gradient-based sample selection step to ensure that only $M$ representative samples are kept for the next GP update.
4.3 Gradient-based Sample Selection BO
The following outlines the GSSBO implementation details and considerations to improve the optimization process, effectively addressing practical challenges. We highlight the key insight of this subsection: we tackle the scalability of BO by maintaining a subset of the most representative and informative samples that are selected based on gradient information.
Detailed Implementations. In the initialization phase, $n_{0}$ initial samples $\{(x_{i},y_{i})\}_{i=1}^{n_{0}}$ are observed; and the initial sample set $D$, the buffer size $M$, and total budget $N$ are specified. In each iteration, the GP posterior is updated on the current sample set $D$, an acquisition function (e.g., UCB) is built to select the next point $x_{t}$, the corresponding observation $y_{t}=f(x_{t})$ is obtained, and the sample $(x_{t},y_{t})$ is added to $D$. To manage each iteration’s computational complexity, a buffer check and gradient-based selection step are performed. Specifically, if the current size of $D$ is less than or equal to $M$, the GP is updated using all samples. Otherwise, a gradient-based sample selection step is performed to identify a set of the most representative samples. Note that the newly acquired sample $(x_{t},y_{t})$ are always added into the subset, as they provide base information for the GP model and ensure that recently observed information is always retained, respectively. Besides the newly observed samples, (M1)(M-1) samples are selected by minimizing the sum of pairwise cosine similarities among their gradients. The resulting subset $\mathcal{U}$, containing $M$ samples, is then used to fit the GP model. The complete procedure is outlined in Algorithm 4.3.
(1) Dynamic Buffer Size. In practice, the buffer size should be prespecified by the users. However, the value is often unavailable in advance. Instead, we propose a dynamic adjustment mechanism to determine the buffer size. We define a tolerable maximum factor $Z$ to accelerate GP computations. Let $\bar{T}$ be the average wall-clock time for a single initial iteration and $T_{\text{current}}$ be the current iteration’s computation time. If $T_{\text{current}}$ exceeds the user-specified threshold $Z\times\bar{T}$, the buffer size is set to be the number of all current samples, i.e., $M=|D|$. This adaptive strategy ensures that the algorithm balances computational efficiency with the goal of utilizing as much data as possible, thereby maintaining high predictive accuracy without incurring excessive costs.
(2) Preserving Latest Observations. During the procedure, the newly acquired sample, $(x_{t},y_{t})$, is also included in the subset. This ensures that the GP model incorporates the latest data, maintaining its relevance and accuracy. Consequently, the algorithm prevents valuable information from being prematurely excluded. Additionally, this essentially alleviates a limitation of sparse GP in BO [McIntire et al., 2016]: the constrained representation size may hinder the full integration of new observations into the model. This will also help to escape local optima from iteratively selecting a subset of “locally optimal” samples, since this observed sample will not be observed in next iteration.
Here we highlight the difference between our method and SparseGP. SparseGP methods require an additional correction vector or term to compensate for the diagonal discrepancies between the low-rank kernel matrix and the high-rank kernel matrix. In contrast, our method leverages gradient information to select the most representative samples, thereby constructing a low-rank approximation that directly approximates the full kernel matrix without the need for extra correction vectors. This not only simplifies the model structure but also reduces the additional computational overhead. The overall speed of GSSBO is often faster than Sparse GP, which requires multiple rounds of optimization for variational inference.
Algorithm 1 Gradient-based Sample Selection BO
1:Initialization: Obtain $n_{0}$ initial samples $D=\{(x_{i},y_{i})\}_{i=1}^{n_{0}}$, and fit an initial GP model. Set buffer size $M>n_{0}$, total budget NN, average initial iteration time $\bar{T}$, and threshold factor $Z$. Initialize switched $\leftarrow$ false.
2:for $t=n_{0}+1$ to NN do
3:  Select $x_{t}=\arg\max_{x}\;\alpha\bigl(x;\,p(f\mid D)\bigr),$ where $\alpha$ is the acquisition function.
4:  Evaluate $y_{t}=f(x_{t})$ and set $D\leftarrow D\cup\{(x_{t},y_{t})\}$.
5:  if not switched then
6:   Let $T_{\text{current}}$ be the current iteration time.
7:   if $T_{\text{current}}>Z\times\bar{T}$ then
[…]
9:   end if
10:  end if
11:  if switched then
12:   Compute gradients $g_{i}$ for all $(x_{i},y_{i})\in D$.
13:   Form subset $\mathcal{U}$ by forcing in $(x_{t},y_{t})$.
14:   Select the remaining M1M-1 samples from $D$ via the gradient-based criterion.
15:   Update the GP using the $M$ samples in $\mathcal{U}$.
16:  else
17:   Update the GP using all samples in $D$.
18:  end if
19:end for
5 Theoretical Analysis
Gaussian Process Upper Confidence Bound (GP-UCB [Srinivas et al., 2009]) is a popular algorithm for sequential decision-making problems. We propose an extension to GP-UCB by incorporating gradient-based sampling. Some assumptions and notations follow  [Srinivas et al., 2009]. Here we analyze the error of the subset fitted GP and prove that the regret of the GSSBO with GP-UCB algorithm is bounded.
Theorem 5.1.
(Error in the Subset-Fitted GP) This theorem establishes bounds on the difference between the posterior mean and variance under a subset fitted GP approximation and those of the full set fitted GP. For a noisy sample $\mathbf{y}=f(\mathbf{X})+\bm{\epsilon},~\bm{\epsilon}\sim\mathcal{N}(0,\sigma_{n}^{2}\mathbf{I}),$i.i.d. Given a GP with kernel matrix $\mathbf{K}_{\mathcal{D}\mathcal{D}}$, a low-rank approximation $\hat{\mathbf{K}}=\mathbf{K}_{\mathcal{D}\mathcal{U}}\mathbf{K}_{\mathcal{U}\mathcal{U}}^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}}$ constructed from $M$ inducing samples and a test sample $\mathbf{x}_{*}$, the posterior predictive mean and variance errors satisfy:
[…]
where $\mathbf{k}_{*\mathcal{D}}\in\mathbb{R}^{N}$ means the covariance vector between the test sample $\mathbf{x}_{*}$ and all training samples in $\mathcal{D}$, $C_{M}=\|({\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I}})^{-1}\|\|({\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I}})^{-1}\|$.
To aid in the theoretical analysis, we make the following assumptions.
Assumption 1. Assume there exist constants $a$, $b$, and $L$ such that the kernel function $k(\mathbf{x},\mathbf{x}^{\prime})$ satisfies a Lipschitz continuity condition, providing confidence bounds on the derivatives of the GP sample paths $f$: $P\left(\sup_{\mathbf{x}\in\mathcal{X}}\left|\frac{\partial f}{\partial x_{j}}\right|>L\right)\leq ae^{-L^{2}/b^{2}}\quad\text{for }j=1,\ldots,d.$ A typical example of such a kernel is the squared exponential kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\sigma^{2}\exp\left(-\frac{\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}}{2l^{2}}\right)$, where $l$ is the length-scale parameter and $\sigma^{2}$ represents the noise variance. Then we propagate the error in Theorem 5.1 through the GP posterior to bound the GSSBO regret.
Theorem 5.2.
(Regret Bound for GSSBO with UCB) Let $\mathcal{X}\subseteq[0,r]^{d}$ be compact and convex, $d\in\mathbb{N},r>0$, let $A=\|\mathbf{k}_{*\mathcal{D}}\|\|\mathbf{y}\|\;C_{M}\,(\lambda_{M+1}+\epsilon_{g}\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2})$, $B_{n}=\|\mathbf{k}_{*\mathcal{D}}\|\;\sqrt{\,C_{M}\;\Bigl(\lambda_{M+1}+\epsilon_{g}\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}\Bigr)}$, where $\lambda_{M+1}$ is the (M+1)(M+1)‐th largest eigenvalue of $\mathbf{K}_{\mathcal{D}\mathcal{D}}$, $\epsilon_{g}$ is the Nyström approximation error parameter for gradient based sample selection, and $\sigma_{\min}>0$ satisfies $\sigma(x)\geq\sigma_{\min}$ for all $x\in\mathcal{X}$, where $\sigma(x)$ is posterior standard deviation. Under Assumption 1, for any arbitrarily small $\delta\in(0,1)$, choose $\beta_{n}\;=\;\frac{\sigma_{\min}(n)\,\Bigl[2\log\frac{4\pi_{n}}{\delta}+2d\log\left(n^{2}brd\sqrt{\log\left(\frac{4da}{\delta}\right)}\right)\Bigr]\;-\;A_{n}}{\sigma_{\min}(n)\;+\;B_{n}}\,.$ Where $\sum_{n\geq 1}\pi_{n}^{-1}=1,\pi_{n}>0$. As $n\to\infty$, we obtain a regret bound of $\mathcal{O}^{*}\left(\sqrt{dN\gamma_{T_{N}}}\right)$. Where $\gamma_{T_{N}}$ is the information gain. Specifically, with $C_{1}=\frac{8}{\log(1+\sigma^{-2})},$ we have:
$P\left(R_{N}\leq\sqrt{C_{1}{T_{N}}\beta_{{T_{N}}}\gamma_{T_{N}}}\right)\geq 1-\delta.$ (5)
Theorem 5.3.
(Two-Phase Regret) Let $T_{N}$} be the total number of rounds. In the first $T_{M}$} rounds, one applies the full GP-UCB. Subsequently, from round $T_{M+1}$} to $T_{N}$}, one switches to the gradient-based subset strategy. The total regret satisfies $R_{T_{N}}\;=\;R_{{T_{M}}}^{(\text{full})}\;+\;R_{T_{N-M}}^{(\text{selected})},$ where $R_{{T_{M}}}^{(\text{full})}\leq\sqrt{\,C_{1}\,{T_{M}}\,\beta_{T_{M}}\,\gamma_{T_{M}}}+2$ and $R_{T_{N-M}}^{(\text{selected})}\leq\sqrt{\,C_{1}\,(T_{N}-T_{M})\,\beta_{T_{N-M}}\,\gamma_{T_{N-M}}}$.
The sketch proof for the main theorem is relegated to the Appendix. The main theoretical challenge lies in evaluating the error between the low-rank approximation $\hat{\mathbf{K}}$ and the full $\mathbf{K}_{\mathcal{D}\mathcal{D}}$. By invoking spectral norm inequalities and the Nyström approximation theory, $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|$ can be bounded. We prove that select $\mathcal{U}$ by maximizing gradient‐direction diversity coincides with the greedy Nyström column‐selection, then derive select $\mathcal{U}$ by maximizing gradient‐direction diversity is better than random selection in standard Nyström approximation. Then, we merge the resulting linear and $\beta$-scaled error terms into a single penalty in the UCB construction. We also determine the smallest subset size $M_{\min}$. This establishes a regret bound for GSSBO, which is similar to that of classical GP-UCB. From a practical relevance perspective, Theorem 5.1 indicates that limiting the GP to a smaller, well-chosen subset does not substantially degrade posterior accuracy in either the mean or variance estimates. Restricting the subset size $M$ confers significant computational savings while ensuring performance closely matches that of a standard GP-UCB using all samples. Of note, we also observe that, compared with the classical UCB results, our GSSBO retains the same fundamental structure of an upper confidence bound approach. Still, it restricts the GP fitting to a gradient-based sample subset, lowering computational costs.
6 Experiments
In this section, we conduct numerical experiments to illustrate the superior efficiency of GSSBO. The objective of the numerical experiments is threefold: (1) to evaluate computational efficiency; (2) to assess optimization performance; and (3) to validate the theoretical analysis. To assess the performance of our proposed methods, we test five benchmark functions, Eggholder2, Hart6, Levy20, Powell50, Rastrigin100 and Neural Architecture Search (NAS).
Figure 2: Cumulative regret of algorithms on synthetic and real-world test problem experiments.
Figure 3: Cumulative time cost of algorithms.
Figure 4: Sensitivity analysis of $Z$.
Experimental Setup.
We choose UCB as the acquisition function in GSSBO, and compare GSSBO with the following benchmarks: (1) Standard GP-UCB [Srinivas et al., 2009]; (2) Random Sample Selection GP-UCB (RSSBO), which mirrors our approach in restricting the sample set size but chooses which samples randomly; (3) VecchiaBO [Jimenez and Katzfuss, 2023], which utilizes the Vecchia approximation method in BO; (4) LR-First m [Williams and Seeger, 2000], which uses a low-rank approximation based on the first $m$ samples. (5) SparseGP [Lawrence et al., 2002], which introduces a small number of inducing points to obtain a sparse approximation of GP; (6) IPA [Moss et al., 2023], which multiplies an expected‐improvement‐based quality function by the GP posterior and selects inducing points regarding quality‐diversity. and (7) mini-META [Calandriello et al., 2022] scaling GP optimization by repeatedly evaluating each selected point until its posterior uncertainty falls below a preset threshold. (More baseline and high dimensional task can be found in the appendix.) We employ a Matérn $5/2$ kernel for the GP, with hyperparameters learned via maximum likelihood estimation. Both GSSBO and RSSBO use the same buffer size $M$, dynamically adjusted by a parameter $Z=4$, $M$ is usually around 100 and will change with $Z$. The gradient-perturbation noise is set to $\sigma^{2}=0.01$. The size of the initial set is 20. Each experiment is repeated 50 times, and the total number of iterations is 1000. All experiments were conducted on a MacBook Pro with Apple M2 Pro (10-core CPU, 16 GB unified RAM).
6.1 Synthetic Test Problems
Computational Efficiency Analysis. Figure 3 compares the cumulative runtime(in seconds) over $1000$ iterations on low-dimensional Hart6 and high-dimensional Powell50 (results for other test functions are provided in the appendix due to space constraints). In both plots, VecchiaBO incurs a rapidly accelerating runtime, whereas GSSBO and RSSBO remain notably lower than GP-UCB. While VecchiaBO reduces the cost of GP fitting by conditioning on nearest neighbors, its runtime is dominated by the costly maintenance of a structured neighbor graph, which scales poorly with dimensionality and sample size. The running speeds of LR-First, SparseGP, and mini-META are all improved compared to the standard GP-UCB. IPA has no obvious advantage in terms of time consumption because it needs to calculate the sample quality. In contrast, GSSBO and RSSBO cost much less time. Because we use an efficient pilot-based method to compute gradients with minimal computational overhead, the fitting cost per iteration of GSSBO remains around $\mathcal{O}(M^{3})$ once we restrict the active subset to size $M\ll n$. The GSSBO and the random one are often similar in runtime, though the GSSBO can be slightly higher due to the overhead of the gradient-based sample selection. By iteration 1000, the total running time of GSSBO on both functions is only about $10\%$ of that of standard GP-UCB. This advantage of GSSBO becomes more pronounced as $n$ increases.
Optimization Performance Analysis. Figure 2 compares methods on multiple functions, evaluating cumulative regret. Overall, GSSBO achieves comparable performance with the Standard GP-UCB while outperforming the RSSBO. In low-dimensional problems such as Eggholder2 and Hart6, GSSBO has a subtle gap with Standard GP-UCB and VecchiaBO. In contrast, GSSBO significantly outperforms other methods. In high-dimensional settings, such as on Levy20, Powell50, and Rastrigin100, GSSBO achieves the smallest cumulative regret, surpassing other methods. From the experimental results, we can observe that the cumulative regret of our algorithm is sublinear, which is consistent with the theoretical results. In particular, GSSBO achieves these results with a significant reduction in computation time, as shown in Figure 3. GSSBO strikes a combination of performance and efficiency in scalable optimization tasks by maintaining near-baseline regret while significantly improving computational efficiency. The reason GP‑UCB outperforms most other baselines in Figure 2 is that it employs full-data standard GP updates, whereas the other baselines use approximate algorithms that incur information loss, leading to their inferior performance.
Sensitivity analysis of Hyperparameter $Z$. We further examine how the dynamic buffer parameter $Z$ affects GSSBO. Figure 4 presents results on two functions: Hart6 and Powell50. For RSSBO, $Z$ remains fixed at 4, whereas for GSSBO, we vary $Z\in\{2,4,6,8\}$. On Hart6, larger $Z$ consistently boosts the GSSBO’s performance toward that of Standard GP-UCB, while the RSSBO lags in cumulative regret. Intuitively, for low-dimensional problems, allowing the model to retain more samples helps preserve important information, bridging the gap with the Standard GP-UCB baseline. In contrast, in Powell50, smaller $Z$ leads to slightly better performances for GSSBO, reflecting the benefit of subset updates in high-dimensional landscapes. In general, low-dimensional tasks benefit from a larger $Z$, while high-dimensional problems perform better with more aggressive subset limiting, $Z$ can be effectively tuned to match the complexity of tasks.
6.2 Real-World Application
To assess the applicability of GSSBO in real-world applications, we use a diabetes-detection problem from the UCI repository [Dua and Graff, 2017]. We modeled the problem of searching for optimal hyperparameters as a BO problem. Specifically, each query $(x_{t},y_{t})$ corresponds to a choice of ([32,128],[1e6,1.0],[1e6,1.0],[1,8][32,128],[1e-6,1.0],[1e-6,1.0],[1,8] for batch size, Learning Rate, Learning Rate decay, hidden dim), where $y_{t}$ is the test classification error. The results in Figure 2 indicate that GSSBO outperforms all competitors.
7 Conclusion
BO is known to be effective for optimization in settings where the objective function is expensive to evaluate. In large-budget scenarios, the use of a full GP model can slow the convergence of BO, leading to poor scaling in these cases. In this paper, we investigated the use of gradient-based sample selection to accelerate BO, demonstrating how a carefully constructed subset, guided by gradient information, can serve as an efficient surrogate for the full sample set, significantly enhancing the efficiency of the BO process. As we have shown in a comprehensive set of experiments, the proposed GSSBO shows its ability to significantly reduce computational time while maintaining competitive optimization performance. Synthetic and real-world benchmarks highlight its scalability and practical utility across various problem dimensions. The sensitivity analysis further showcases the adaptability of the method to different parameter settings. Overall, these findings underline the potential of gradient-based sample selection in addressing the BO scaling challenges.
Appendix
Appendix A Appendix Theoretical Analysis
A.1 Theorem 1:Analysis on subset GP:
Consider a GP model, where we assume $f(\mathbf{x})\sim\mathcal{GP}(0,k(\mathbf{x},\mathbf{x}^{\prime})),$ and given samples $\mathcal{D}=\{(\mathbf{x}_{1},y_{1}),\ldots,(\mathbf{x}_{N},y_{N})\}$, where $\mathbf{x}_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\mathbb{R}$, we have a noise model: $\mathbf{y}=f(\mathbf{X})+\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(0,\sigma_{n}^{2}\mathbf{I}),$ where $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{N}]^{\top}\in\mathbb{R}^{N\times d}$ and $\mathbf{y}=[y_{1},y_{2},\ldots,y_{N}]^{\top}\in\mathbb{R}^{N}$. The posterior predictive distribution for a test point $\mathbf{x}_{*}$ is Gaussian with the following mean and variance: $\mu(\mathbf{x}_{*})=\mathbf{k}_{*}^{\top}(\mathbf{K}+\sigma_{n}^{2}\mathbf{I})^{-1}\mathbf{y},$ $\sigma^{2}(\mathbf{x}_{*})=k_{**}-\mathbf{k}_{*}^{\top}(\mathbf{K}+\sigma_{n}^{2}\mathbf{I})^{-1}\mathbf{k}_{*},$ where $\mathbf{K}$ is the $N\times N$ kernel matrix evaluated at the training samples, i.e., $\mathbf{K}_{ij}=k(\mathbf{x}_{i},\mathbf{x}_{j})$; $\mathbf{k}_{*}\in\mathbb{R}^{N}$ is the vector of covariances between the test point $\mathbf{x}_{*}$ and all training points, i.e., $(\mathbf{k}_{*})_{i}=k(\mathbf{x}_{*},\mathbf{x}_{i})$; $k_{**}=k(\mathbf{x}_{*},\mathbf{x}_{*})$ is the kernel evaluated at the test point itself.
Instead of using all $N$ training samples, consider a subset samples $\mathcal{U}=\{\mathbf{u}_{1},\dots,\mathbf{u}_{M}\}$, where $M\ll N$. Define: $\mathbf{K}_{\mathcal{U}\mathcal{U}}\in\mathbb{R}^{M\times M},\quad\mathbf{K}_{\mathcal{D}\mathcal{U}}\in\mathbb{R}^{N\times M},\quad\mathbf{K}_{\mathcal{U}\mathcal{D}}=\mathbf{K}_{\mathcal{D}\mathcal{U}}^{\top},$ where $\mathbf{K}_{\mathcal{U}\mathcal{U}}$ is the kernel matrix among the $M$ inducing points, and $\mathbf{K}_{\mathcal{D}\mathcal{U}}$ represents the covariances between the full training samples in $\mathcal{D}$ and the inducing points in $\mathcal{U}$. Using Subset of Regressors (SoR), a low-rank approximation to the kernel matrix $\mathbf{K}_{\mathcal{D}\mathcal{D}}\in\mathbb{R}^{N\times N}$ is given by: $\hat{\mathbf{K}}=\mathbf{K}_{\mathcal{D}\mathcal{U}}\mathbf{K}_{\mathcal{U}\mathcal{U}}^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}}.$
Error Characterization by Kernel Approximation and Variance Error Analysis
We aim to bound the difference between the exact posterior distribution and the approximate posterior distribution. The differences in the posterior predictive mean and variance can be expressed as: $\Delta\mu(\mathbf{x}_{*})=\mu(\mathbf{x}_{*})-\tilde{\mu}(\mathbf{x}_{*}),$ $\Delta\sigma^{2}(\mathbf{x}_{*})=\sigma^{2}(\mathbf{x}_{*})-\tilde{\sigma}^{2}(\mathbf{x}_{*}).$ Starting with the mean difference, the exact posterior predictive mean and the approximate mean are given by: $\mu(\mathbf{x}_{*})=\mathbf{k}_{*\mathcal{D}}^{\top}(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}\mathbf{y},$ $\tilde{\mu}(\mathbf{x}_{*})=\mathbf{k}_{*\mathcal{D}}^{\top}(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}\mathbf{y}.$ Thus, the mean difference becomes:
$\Delta\mu(\mathbf{x}_{*})=\mathbf{k}_{*\mathcal{D}}^{\top}\left[(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}-(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}\right]\mathbf{y}.$ (6)
[…]
The approximate posterior predictive variance is given by: $\tilde{\sigma}^{2}(\mathbf{x}_{*})=k_{**}-\mathbf{k}_{*\mathcal{D}}^{\top}(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}\mathbf{k}_{*\mathcal{D}},$ where $\hat{\mathbf{K}}=\mathbf{K}_{\mathcal{D}\mathcal{U}}\mathbf{K}_{\mathcal{U}\mathcal{U}}^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}}$ is the low-rank approximation to $\mathbf{K}_{\mathcal{D}\mathcal{D}}$. Define the variance error as: $\Delta\sigma^{2}(\mathbf{x}_{*})=\sigma^{2}(\mathbf{x}_{*})-\tilde{\sigma}^{2}(\mathbf{x}_{*}).$ The variance error can be expressed as:
$\begin{split}\Delta\sigma^{2}(\mathbf{x}_{*})&=\bigl[\,k_{**}-\mathbf{k}_{*\mathcal{D}}^{\top}\bigl(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I}\bigr)^{-1}\mathbf{k}_{*\mathcal{D}}\bigr]-\bigl[\,k_{**}-\mathbf{k}_{*\mathcal{D}}^{\top}\bigl(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I}\bigr)^{-1}\mathbf{k}_{*\mathcal{D}}\bigr]\\ &=\mathbf{k}_{*\mathcal{D}}^{\top}\bigl[\,(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}-(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}\bigr]\mathbf{k}_{*\mathcal{D}}\,.\end{split}$ (8)
From the mean error analysis, we know: $\|(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}-(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}\|\leq C_{M}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|$. Using the spectral norm, the variance error can be bounded as: $|\Delta\sigma^{2}(\mathbf{x}_{*})|=\left|\mathbf{k}_{*\mathcal{D}}^{\top}\left[(\hat{\mathbf{K}}+\sigma_{n}^{2}\mathbf{I})^{-1}-(\mathbf{K}_{\mathcal{D}\mathcal{D}}+\sigma_{n}^{2}\mathbf{I})^{-1}\right]\mathbf{k}_{*\mathcal{D}}\right|.$ Applying the properties of matrix norms and the Cauchy–Schwarz inequality:
[…]
Impact of the “Maximum Gradient Variance” Principle in Selecting Subsamples
The “Maximum Gradient Variance” principle for selecting the inducing points aims to pick points that best capture the main gradient variations in the sample distribution. As $M$ increases and the subset points are chosen more effectively, the approximation $\hat{\mathbf{K}}$ to $\mathbf{K}_{\mathcal{D}\mathcal{D}}$ improves, hence $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|$ decreases. Consider a set of $N$ samples $\mathcal{D}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$ and a corresponding kernel matrix: $\mathbf{K}_{\mathcal{D}\mathcal{D}}\in\mathbb{R}^{N\times N},\quad(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ij}=k(\mathbf{x}_{i},\mathbf{x}_{j}),$ where $k$ is a positive definite kernel. Suppose we have observations $\mathbf{y}\in\mathbb{R}^{N}$ associated with these samples, and a probabilistic model (e.g., a GP) with parameters $\theta$} and the mean function $\mu$. The joint distribution of $\mathbf{y}$ given $\mathcal{D}$ and $\theta$} is: $\mathbf{y}\mid\mathcal{D},{\theta}\sim\mathcal{N}(\mu,\mathbf{K}_{\mathcal{D}\mathcal{D}}),$ where $\mathbf{K}_{\mathcal{D}\mathcal{D}}$ is the covariance matrix induced by the kernel $k$. Define the gradient of the log-posterior (or log-likelihood) with respect to the latent function values $\mathbf{y}$: $g_{i}=\frac{\partial\log p(\mathbf{y}\mid\mathcal{D},{\theta})}{\partial y_{i}}=-(\mathbf{K}_{\mathcal{D}\mathcal{D}}^{-1}(\mathbf{y}-\mu))_{i}.$
Now we force the latest sample $\mathbf{x}_{\mathrm{L}}$ to be included in $\mathcal{U}$. We then choose the remaining M1M-1 samples to maximize gradient variance. We select the M1M-1 samples that maximize the variance of gradient information. We have $\mathcal{U}\;=\;\{\mathbf{x}_{\mathrm{L}}\}\;\cup\;\mathcal{U}^{\prime},$ where $|\mathcal{U}^{\prime}|=M-1$ so that $|\mathcal{U}|=M$, the total subset $\mathcal{U}$ is still of size $M$. We consider a subset $\mathcal{U}\subset\mathcal{D}$ of size $M<N$ to build a low-rank approximation of $\mathbf{K}_{\mathcal{D}\mathcal{D}}$: $\hat{\mathbf{K}}=\mathbf{K}_{\mathcal{D}\mathcal{U}}\mathbf{K}_{\mathcal{U}\mathcal{U}}^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}},$ where $\mathbf{K}_{\mathcal{U}\mathcal{U}}$ and $\mathbf{K}_{\mathcal{D}\mathcal{U}}$ are derived from $\mathcal{U}$. Let $\mathcal{D}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$ and $\mathbf{K}_{\mathcal{D}\mathcal{D}}\in\mathbb{R}^{N\times N}$ be a positive definite kernel matrix with eigen-decomposition: $\mathbf{K}_{\mathcal{D}\mathcal{D}}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{\top},$ where $\bm{\Lambda}=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{N})$ with $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{N}\geq 0$ and $\mathbf{U}=[\mathbf{u}_{1},\ldots,\mathbf{u}_{N}]$. The best rank-$M$ approximation to $\mathbf{K}_{\mathcal{D}\mathcal{D}}$ in spectral norm is: $\mathbf{K}_{\mathcal{D}\mathcal{D}}^{(M)}=\mathbf{U}_{M}\bm{\Lambda}_{M}\mathbf{U}_{M}^{\top},$ where $\mathbf{U}_{M}=[\mathbf{u}_{1},\dots,\mathbf{u}_{M}]$ and $\bm{\Lambda}_{M}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{M})$. By the Eckart–Young–Mirsky theorem: $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\mathbf{K}_{\mathcal{D}\mathcal{D}}^{(M)}\|=\lambda_{M+1}.$ Suppose $\mathcal{U}\subset\mathcal{D}$, $|\mathcal{U}|=M$, produces a approximation: $\hat{\mathbf{K}}=\mathbf{K}_{\mathcal{D}\mathcal{U}}\mathbf{K}_{\mathcal{U}\mathcal{U}}^{-1}\mathbf{K}_{\mathcal{U}\mathcal{D}}.$
[…]
If we choose $\mathcal{U}$ via a greedy procedure that maximizes gradient‐information, then we defines the orthogonal projector onto the selected columns by $\mathbf{P_{\mathcal{U}}}\;=\;\mathbf{K_{D\,U}\,(K_{U\,U})^{-1}K_{U\,D}},$ and measures the subspace approximation error against the true principal eigenspace $\mathrm{span}(U_{M})$ as
$\begin{split}\mathbf{\epsilon_{g}}\;=\;\bigl\|\,(\mathbf{I}-\mathbf{P_{\mathcal{U}}})\,\mathbf{U_{M}}\bigr\|_{F}\;=\;\sqrt{\sum_{k=1}^{M}\bigl\|(\mathbf{I}-\mathbf{P_{\mathcal{U}}})\,\mathbf{u_{k}}\bigr\|^{2}}.\end{split}$ (10)
Then if $\mathcal{U}$ is chosen by Gradient Information, we obtain the analogous Nyström error bound:
$\begin{split}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|\;\leq\;\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\mathbf{K}_{\mathcal{D}\mathcal{D}}^{(M)}\|\;+\;\epsilon_{g}\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}.\end{split}$ (11)
Random select $\mathcal{U}$. If the subset $\mathcal{U}$ is chosen to approximate the principal eigenspace spanned by randomly choose $\mathbf{U}_{M}$, then Nyström approximation theory [Drineas et al., 2005] guarantees that: $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|\leq\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\mathbf{K}_{\mathcal{D}\mathcal{D}}^{(M)}\|+\epsilon\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}=\lambda_{M+1}+\epsilon\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}.$ where $\epsilon$ is an error control parameter related to the number of columns sampled randomly in the approximation. In particular, if one samples $c$ columns uniformly at random , then for any $\delta\in(0,1)$ the following holds with probability at least $1-\delta$ [Drineas et al., 2005]:
$\begin{split}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|\;\leq\;\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\mathbf{K}_{\mathcal{D}\mathcal{D}}^{(M)}\|\;+\;\epsilon\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2},\end{split}$ (12)
where $\epsilon\;=\;\sqrt{\frac{8}{c}\,\ln\frac{2M}{\delta}}\;=\;O\!\Bigl(\tfrac{1}{\sqrt{c}}\Bigr).$
Comparison. Let $\mathbf{K}_{\mathcal{D}\mathcal{D}}\in\mathbb{R}^{N\times N}$ have eigenpairs $(\lambda_{k},u_{k})$, and let $U_{M}=[\,u_{1},\dots,u_{M}\,]\in\mathbb{R}^{N\times M}$ be the matrix of its top $M$ eigenvectors. Suppose we build a Nyström approximation by selecting columns greedily in $M$ steps; let $\mathcal{U}_{t}$ be the chosen indices after $t$ steps and $\mathbf{P_{t}}$ the corresponding orthogonal projector onto the span of those columns. Define the residual energy $E_{t}\;=\;\sum_{k=1}^{M}\|(\mathbf{I}-\mathbf{P_{t}})\mathbf{u_{k}}\|^{2}.$ Observe $\sum_{i=1}^{N}\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{\phi_{i}}\|^{2}=\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{K}\|_{F}^{2}\;\geq\;\|(\mathbf{I}-\mathbf{P_{t}})\,U_{M}\Lambda_{M}^{1/2}\|_{F}^{2}\;\geq\;\lambda_{1}\,\|(\mathbf{I}-\mathbf{P_{t}})\,U_{M}\|_{F}^{2}=E_{t},$ where $\mathbf{\phi_{i}}=\mathbf{K}\,\mathbf{e_{i}}$ and $\Lambda_{M}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{M})$. Hence among the remaining NtN-t columns there is some index $j\notin\mathcal{U}_{t}$ with $\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{\phi_{j}}\|^{2}\;\geq\;\frac{1}{N-t}\sum_{i\notin\mathcal{U}_{t}}\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{\phi_{i}}\|^{2}\;\geq\;\frac{E_{t}}{N-t}.$ Since the greedy rule picks precisely the column that maximizes $\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{\phi_{i}}\|$, the new residual satisfies $E_{t+1}=E_{t}\;-\;\|(\mathbf{I}-\mathbf{P_{t}})\,\mathbf{\phi_{j}}\|^{2}\;\leq\;E_{t}\Bigl(1-\tfrac{1}{N-t}\Bigr).$ Iterating the one‐step bound from $t=0$ to t=M1t=M-1 gives $E_{M}\;\leq\;E_{0}\prod_{t=0}^{M-1}\Bigl(1-\tfrac{1}{N-t}\Bigr)=\;M\;\prod_{k=N-M+1}^{N}\Bigl(1-\tfrac{1}{k}\Bigr),$ since $E_{0}=\|U_{M}\|_{F}^{2}=M$. Using $1-\tfrac{1}{k}\leq e^{-1/k}$, we obtain $E_{M}\;\leq\;M\;\exp\!\Bigl(-\sum_{k=N-M+1}^{N}\tfrac{1}{k}\Bigr)\;\leq\;M\;\exp\!\bigl(-\tfrac{M}{N}\bigr).$ Thus the final projection error $\epsilon_{g}=\sqrt{E_{M}}\;\leq\;\sqrt{M}\,\exp\!\bigl(-\tfrac{M}{2N}\bigr)=O\!\bigl(e^{-\tfrac{M}{2N}}\bigr)$. Because $\epsilon_{g}$ decays exponentially in $M$, it is asymptotically much smaller than the $O(1/\sqrt{M})$ decay of the random‐sampling parameter $\epsilon$. Hence, greedy (gradient‐based) selection yields a Nyström approximation error that is provably tighter than uniform random sampling for the same $M$.
The key insight is that each greedy step removes at least the average of the remaining projection error (or “energy”). If at step $t$ the total unexplained energy is $E_{t}$, then among the NtN-t remaining columns there must be one whose residual contributes at least $E_{t}/(N-t)$. By selecting that column, the algorithm shrinks the leftover energy by a factor of $\bigl(1-\tfrac{1}{N-t}\bigr)$. Repeating this multiplicative reduction $M$ times drives the energy down by a product of $\bigl(1-\tfrac{1}{k}\bigr)$ terms, which is bounded by an exponential $e^{-M/N}$. Hence the projection error $\epsilon_{g}$ decays exponentially in $M$.
Our theoretical analysis dictates that the low-rank Nyström approximation error remains within a small relative tolerance of the full kernel. Formally, we require $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|=\lambda_{M+1}+\epsilon_{g}\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}\leq 0.05\|\mathbf{K}_{\mathcal{D}\mathcal{D}}\|.$ Here, $\lambda_{M+1}$ denotes the $(M+1)$-th eigenvalue of $\mathbf{K}_{\mathcal{D}\mathcal{D}}$, and $\epsilon_{g}$ captures the projection error induced by the greedy column selection. We can solve for the smallest subset size $M_{\min}$ that satisfies this inequality.
Error Bound for UCB under Sparse GP Approximation
Let the UCB for the full GP model be defined as: $\mathrm{UCB}(\mathbf{x}_{*})=\mu(\mathbf{x}_{*})+\beta_{n}\sigma(\mathbf{x}_{*}),$ where $\mu(\mathbf{x}_{*})$ and $\sigma(\mathbf{x}_{*})$ are the posterior predictive mean and standard deviation under the full GP, respectively. For the sparse GP approximation, the UCB is given by: $\tilde{\mathrm{UCB}}(\mathbf{x}_{*})=\tilde{\mu}(\mathbf{x}_{*})+\beta_{n}\tilde{\sigma}(\mathbf{x}_{*}),$ where $\tilde{\mu}(\mathbf{x}_{*})$ and $\tilde{\sigma}(\mathbf{x}_{*})$ are the posterior predictive mean and standard deviation under the sparse GP approximation.
We aim to bound the error: $|\mathrm{UCB}(\mathbf{x}_{*})-\tilde{\mathrm{UCB}}(\mathbf{x}_{*})|,$ in terms of the kernel matrix approximation error $\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|$. Given the bounds:
[…]
The error of UCB in one iteration can be expressed as:
$\bigl|\mathrm{UCB}(\mathbf{x}_{*})-\tilde{\mathrm{UCB}}(\mathbf{x}_{*})\bigr|\;=\;\bigl|\Delta\mu(\mathbf{x}_{*})+\beta_{n}\,\Delta\sigma(\mathbf{x}_{*})\bigr|\;\;\leq\;\;\bigl|\Delta\mu(\mathbf{x}_{*})\bigr|\;+\;\beta_{n}\,\bigl|\Delta\sigma(\mathbf{x}_{*})\bigr|.$ (14)
Substituting the bounds for $|\Delta\mu(\mathbf{x}_{*})|$ and $|\Delta\sigma(\mathbf{x}_{*})|$, we have:
$|\mathrm{UCB}(\mathbf{x}_{*})-\tilde{\mathrm{UCB}}(\mathbf{x}_{*})|\leq\|\mathbf{k}_{*\mathcal{D}}\|\|\mathbf{y}\|C_{M}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|+\beta_{n}\|\mathbf{k}_{*\mathcal{D}}\|\sqrt{C_{M}\|\mathbf{K}_{\mathcal{D}\mathcal{D}}-\hat{\mathbf{K}}\|}$. (15)
Using the Nyström approximation error bound $\|\mathbf{K}-\hat{\mathbf{K}}\|\;\leq\;\lambda_{M+1}+\epsilon_{g}\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2},$ the UCB error can be further bounded as:
$\begin{split}\bigl|\mathrm{UCB}(\mathbf{x}_{*})-\tilde{\mathrm{UCB}}(\mathbf{x}_{*})\bigr|&\leq\|\mathbf{k}_{*\mathcal{D}}\|\,\|\mathbf{y}\|\,C_{M}\Bigl(\lambda_{M+1}+\epsilon_{g}\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}\Bigr)\\ &\quad+\beta_{n}\|\mathbf{k}_{*\mathcal{D}}\|\,\sqrt{\,C_{M}\Bigl(\lambda_{M+1}+\epsilon_{g}\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}\Bigr)}\,.\end{split}$ (16)
Merging Linear and $\beta$-Proportional Terms into a Single Penalty
We consider a GP scenario where the full GP-UCB at a point $\mathbf{x}$ is given by $\mathrm{UCB}(\mathbf{x})=\;\mu(\mathbf{x})+\;\beta_{n}\,\sigma(\mathbf{x}),$ while its approximate counterpart is $\tilde{\mathrm{UCB}}(\mathbf{x})=\;\tilde{\mu}(\mathbf{x})+\;\beta_{n}\,\tilde{\sigma}(\mathbf{x}).$ We have established the following pointwise error bound: $\bigl|\mathrm{UCB}(\mathbf{x})-\tilde{\mathrm{UCB}}(\mathbf{x})\bigr|\;\leq\;A_{n}+\beta_{n}\,B_{n},$ where
$A_{n}=\|\mathbf{k}_{*\mathcal{D}}\|\;\|\mathbf{y}\|\;C_{M}\;\Bigl(\lambda_{M+1}+\epsilon_{g}\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}\Bigr),~B_{n}=\|\mathbf{k}_{*\mathcal{D}}\|\;\sqrt{\,C_{M}\;\Bigl(\lambda_{M+1}+\epsilon_{g}\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}\Bigr)}.$ (17)
Here, $A_{n},B_{n}$ are determined by the number of iterations $n$, the same below. We now absorb both terms into a single, nn–dependent penalty. Define $\sigma_{\min}(n)\;=\;\min_{x\in\mathcal{X}}\,{\sigma}_{n}(x),\qquad\delta_{n}\;=\;\frac{A_{n}+\beta_{n}\,B_{n}}{\sigma_{\min}(n)},\qquad\tilde{\beta}_{n}\;=\;\beta_{n}+\delta_{n}.$ Since for all $x$, ${\sigma}_{n}(x)\geq\sigma_{\min}(n)>0$, it follows that $A_{n}+\beta_{n}\,B_{n}=\;\delta_{n}\,\sigma_{\min}(n)\;\leq\;\delta_{n}\,{\sigma}_{n}(x),$ and hence
$\bigl|\mathrm{UCB}(\mathbf{x})-\tilde{\mathrm{UCB}}(\mathbf{x})\bigr|\;\leq\;A_{n}+\beta_{n}\,B_{n}\;\leq\;\delta_{n}\,{\sigma}_{n}(\mathbf{x})=\frac{A_{n}+\beta_{n}\,B_{n}}{\sigma_{\min}(n)}{\sigma}_{n}(\mathbf{x}).$ (18)
Therefore, we may write the unified approximate UCB as
$\tilde{\mathrm{UCB}}(\mathbf{x})={\mu}(\mathbf{x})+\beta_{n}\,{\sigma}_{n}(\mathbf{x})+\delta_{n}\,{\sigma}_{n}(x)=\;{\mu}(\mathbf{x})+\;\tilde{\beta}_{n}\,{\sigma}_{n}(\mathbf{x}),$ (19)
where $\tilde{\beta}_{n}=\;\beta_{n}+\;\frac{A_{n}+\beta_{n}\,B_{n}}{\sigma_{\min}(n)}.$ This completes the single‐penalty construction.
A.2 Theorem 2:Analysis on Regret Bound:
GP-UCB is a popular algorithm for sequential decision-making problems. We propose an extension to GP-UCB by incorporating gradient-based sampling. In this section, we prove that the regret of the GP-UCB algorithm with gradient-based sampling is bounded. We show that by selecting the subset of samples with the highest variance, we can achieve a regret bound. This approach leverages the information gained from gradient-based sampling to provide a robust regret bound. To aid in the theoretical analysis, we make the following assumptions.
Assumption 1: Assume there exist constants $a$, $b$, and $L$ such that the kernel function $k(\mathbf{x},\mathbf{x}^{\prime})$ satisfies a Lipschitz continuity condition, providing confidence bounds on the derivatives of the GP sample paths $f$:
$P\left(\sup_{\mathbf{x}\in\mathcal{X}}\left|\frac{\partial f}{\partial x_{j}}\right|>L\right)\leq ae^{-L^{2}/b^{2}}\quad\text{for }j=1,\ldots,d.$ (20)
A typical example of such a kernel is the squared exponential kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\sigma^{2}\exp\left(-\frac{\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}}{2l^{2}}\right)$, where $l$ is the length-scale parameter and $\sigma^{2}$ represents the noise variance. This condition aligns with standard assumptions in the regret analysis of BO, as detailed by Srinivas et al. (2010). We now present the main theorem on the cumulative regret bound for the GSSBO.
$\tilde{\mathrm{UCB}}(\mathbf{x})\;=\;\mu(\mathbf{x})\;+\;(\beta_{n}+\frac{A_{n}+\beta_{n}\,B_{n}}{\sigma_{\min}(n)})\,\sigma(\mathbf{x}).$ (21)
Theorem A.1.
Let $\mathcal{X}\subset[0,r]^{d}$ be compact and convex, $d\in\mathbb{N}$, $r>0$. Under Assumption 1, for any arbitrarily small $\delta\in(0,1)$, choose $\tilde{\beta_{n}}=2\log\frac{4\pi_{n}}{\delta}+2d\log\left(n^{2}brd\sqrt{\log\left(\frac{4da}{\delta}\right)}\right),$i.e.,
$\beta_{n}\;=\;\frac{\sigma_{\min}(n)\,\Bigl[2\ln\!\frac{4\pi_{n}}{\delta}\;+\;2d\,\ln\!\Bigl(n^{2}\,b\,r\,d\;\sqrt{\ln\!\tfrac{4da}{\delta}}\Bigr)\Bigr]\;-\;A_{n}}{\sigma_{\min}(n)\;+\;B_{n}}\,.$ (22)
where $\sum_{n\geq 1}\pi_{n}^{-1}=1,\quad\pi_{n}>0$. As $n\to\infty$, we obtain a regret bound of $\mathcal{O}^{*}(\sqrt{d{T_{N}}\gamma_{T_{N}}})$. Specifically, with $C_{1}=\frac{8}{\log(1+\sigma^{-2})},$ we have:
$P\left(R_{T_{N}}\leq\sqrt{C_{1}{T_{N}}\beta_{{T_{N}}}\gamma_{T_{N}}}\right)\geq 1-\delta.$ (23)
Lemma A.2.
For any arbitrarily small $\delta_{1}\in(0,1)$, choose $\tilde{\beta_{n}}=2\log\frac{\pi_{n}}{\delta_{1}}$, i.e., $\beta_{n}\;=\;\frac{\sigma_{\min}(n)\,(2\log\frac{\pi_{n}}{\delta_{1}})\;-\;A_{n}}{\sigma_{\min}(n)\;+\;B_{n}}\,.$ , where $\sum_{n\geq 1}\pi_{n}^{-1}=1$, $\pi_{n}>0$, then we have
$P\left(|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})|\leq d\tilde{\beta_{n}}^{1/2}\sigma_{n-1}(\mathbf{x}_{n})\right)\geq 1-\delta$ (24)
Proof.
Assuming we are at stage $n$, all past decisions $\mathbf{x}_{1:n-1}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n-1}\}$ made after the initial design are deterministic given $\mathbf{y}_{1:n-1}=\{y_{1},\ldots,y_{n-1}\}$. For any $\mathbf{x}_{n}\in\mathbb{R}^{d}$, we have $f(\mathbf{x}_{n})\sim\mathcal{N}(\mu_{n-1}(\mathbf{x}_{n}),\sigma^{2}_{n-1}(\mathbf{x}_{n}))$. For a standard normal variable $r\sim\mathcal{N}(0,1)$, the probability of being above a certain constant $c$ is written as:
$\begin{split}P(r>c)&=\frac{1}{\sqrt{2\pi}}\int_{c}^{\infty}e^{-r^{2}/2}\,dr\\[6.0pt] &=e^{-c^{2}/2}\,\frac{1}{\sqrt{2\pi}}\int_{c}^{\infty}e^{-\frac{(r-c)^{2}}{2}-c(r-c)}\,dr\\[6.0pt] &\leq e^{-c^{2}/2}\,\frac{1}{\sqrt{2\pi}}\int_{c}^{\infty}e^{-\frac{(r-c)^{2}}{2}}\,dr\\[6.0pt] &=e^{-c^{2}/2}\,P(r>0)\\[6.0pt] &=\tfrac{1}{2}\,e^{-c^{2}/2}\,.\end{split}$ (25)
where the inequality holds due to the fact that $e^{-c(r-c)}\leq 1$ for $r\geq c>0$. Plugging in $r=\frac{f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})}{\sigma_{n-1}(\mathbf{x}_{n})}$ and $c=\tilde{\beta_{n}}^{1/2}$, we have:
$P\left(\left|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})\right|>\tilde{\beta_{n}}^{1/2}\sigma_{n-1}(\mathbf{x}_{n})\right)\leq e^{-\frac{\tilde{\beta_{n}}}{2}}.$ (26)
Equivalently,
$P\left(\left|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})\right|\leq\tilde{\beta_{n}}^{1/2}\sigma_{n-1}(\mathbf{x}_{n})\right)\geq 1-e^{-\frac{\tilde{\beta_{n}}}{2}}.$ (27)
Choosing $e^{-\frac{\tilde{\beta_{n}}}{2}}=\frac{\delta}{\pi_{n}}$, i.e., $\tilde{\beta_{n}}=2\log\frac{\pi_{n}}{\delta}$, and applying the union bound for all possible values of stage $n$, we have:
$P\left(\left|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})\right|\leq\tilde{\beta_{n}}^{1/2}\sigma_{n-1}(\mathbf{x}_{n})\right)\geq 1-\sum_{n\geq 1}\frac{\delta}{\pi_{n}}=1-\delta.$ (28)
where we have used the condition that $\sum_{n\geq 1}\pi_{n}^{-1}=1$, which can be obtained by setting $\pi_{n}=\frac{\pi^{2}n^{2}}{6}$.
To facilitate the analysis, we adopt a stage-wise discretization $\mathcal{X}_{n}\subset\mathcal{X}$, which is used to obtain a bound on $f(\mathbf{x}^{*})$.
Lemma A.3.
For any arbitrarily small $\delta\in(0,1)$, choose $\tilde{\beta_{n}}=2\log\frac{|\mathcal{X}_{n}|\pi_{n}}{\delta},i.e.,$ $\beta_{n}\;=\;\frac{\sigma_{\min}(n)\,(2\log\frac{|\mathcal{X}_{n}|\pi_{n}}{\delta})\;-\;A_{n}}{\sigma_{\min}(n)\;+\;B_{n}}\,.$ where $\sum_{n\geq 1}\pi_{n}^{-1}=1,\quad\pi_{n}>0,$ then we have
$P\left(\left|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})\right|\leq\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n})\right)\geq 1-\delta\quad\text{for }\forall\mathbf{x}_{n}\in\mathcal{X}_{n},\forall n\geq 1.$ (29)
Proof.
Based on Lemma 4.2, we have that for each $\mathbf{x}_{n}\in\mathcal{X}_{n}$,
$P\left(\left|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})\right|\leq\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n})\right)\geq 1-e^{-\frac{\tilde{\beta_{n}}}{2}}.$ (30)
Applying the union bound gives:
$P\left(\left|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})\right|\leq\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n})\right)\geq 1-|\mathcal{X}_{n}|e^{-\frac{\tilde{\beta_{n}}}{2}},\quad\forall\mathbf{x}_{n}\in\mathcal{X}_{n}.$ (31)
Choosing $|\mathcal{X}_{n}|e^{-\frac{\tilde{\beta_{n}}}{2}}=\frac{\delta}{\pi_{n}}$, i.e., $\tilde{\beta_{n}}=2\log\frac{|\mathcal{X}_{n}|\pi_{n}}{\delta},$ and applying the union bound for all possible values of stage $n$, we have:
$P\left(\left|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})\right|\leq\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n})\right)\geq 1-\sum_{n\geq 1}\frac{\delta}{\pi_{n}}=1-\delta,$ (32)
where $\sum_{n\geq 1}\pi_{n}^{-1}=1$, $\forall\mathbf{x}_{n}\in\mathcal{X}_{n}$, and $\forall n\geq 1$.
Lemma A.4.
For any arbitrarily small $\delta\in(0,1)$, choose $\tilde{\beta_{n}}=2\log\frac{2\pi_{n}}{\delta}+2d\log\left(n^{2}brd\sqrt{\log\left(\frac{2da}{\delta}\right)}\right),i.e.,$ $\beta_{n}\;=\;\frac{\sigma_{\min}(n)\,\Bigl[2\log\frac{2\pi_{n}}{\delta}+2d\log\left(n^{2}brd\sqrt{\log\left(\frac{2da}{\delta}\right)}\right)\Bigr]\;-\;A_{n}}{\sigma_{\min}(n)\;+\;B_{n}}\,.$ where $\sum_{n\geq 1}\pi_{n}^{-1}=1$, $\pi_{n}>0$, $d\in\mathbb{N}$ is the dimensionality of the feature space, and $r>0$ is the length of the domain in a compact and convex set $\mathcal{X}\subset[0,r]^{d}$. Given constants $a$, $b$ and $L$, assume that the kernel function $k(\mathbf{x},\mathbf{x}^{\prime})$ satisfies the following Lipschitz continuity for the confidence bound of the derivatives of GP sample paths $f$:
$P\left(\sup_{\mathbf{x}\in\mathcal{X}}\left|\frac{\partial f}{\partial x_{j}}\right|>L\right)\leq ae^{-L^{2}/b^{2}},\quad j=1,\dots,d,$ (33)
then we have
$P\left(\left|f(\mathbf{x}^{*})-\mu_{n-1}([\mathbf{x}^{*}]_{n})\right|\leq d\tilde{\beta_{n}^{\frac{1}{2}}}\sigma_{n-1}([\mathbf{x}^{*}]_{n})+\frac{1}{n^{2}}\right)\geq 1-\delta,\quad\forall n\geq 1.$ (34)
Proof.
For $\forall j$, $\mathbf{x}\in\mathcal{X}$, applying the union bound on the Lipschitz continuity property gives:
$P\left(\sup_{\mathbf{x}\in\mathcal{X}}\left|\frac{\partial f}{\partial x_{j}}\right|<L\right)\geq 1-dae^{-L^{2}/b^{2}}$ (35)
which suggests that:
$P\left(\left|f(\mathbf{x})-f(\mathbf{x}^{\prime})\right|\leq L\|\mathbf{x}-\mathbf{x}^{\prime}\|_{1}\right)\geq 1-dae^{-L^{2}/b^{2}}\quad\forall\mathbf{x}\in\mathcal{X}$ (36)
which is a confidence bound that applies to $\mathbf{x}^{*}$ as well. For a discretization $\mathcal{X}_{n}$ of size $(\tau_{n})^{d}$, i.e., each coordinate space of $\mathcal{X}_{n}$ has a total of $\tau_{n}$ discrete points, we have the following bound on the closest point $\mathbf{x}]_{n}$ to $\mathbf{x}$ in $\mathcal{X}_{n}$ to ensure a dense set of discretizations:
$\|\mathbf{x}-[\mathbf{x}]_{n}\|_{1}\leq\frac{rd}{\tau_{n}}.$ (37)
Now, setting $dae^{-L^{2}/b^{2}}=\frac{\delta}{2}$, i.e., $L=b\sqrt{\log\left(\frac{2da}{\delta}\right)}$, gives the following:
$P\left(\left|f(\mathbf{x})-f(\mathbf{x}^{\prime})\right|\leq b\sqrt{\log\left(\frac{2da}{\delta}\right)}\|\mathbf{x}-\mathbf{x}^{\prime}\|_{1}\right)\geq 1-\frac{\delta}{2}\quad\forall\mathbf{x}\in\mathcal{X}.$ (38)
Thus, switching to the discretized space $\mathcal{X}_{n}$ at any stage $n\in\mathbb{R}$ and choosing $\mathbf{x}^{\prime}=[\mathbf{x}]_{n}$ gives:
$P\left(\left|f(\mathbf{x})-f([\mathbf{x}]_{n})\right|\leq brd\sqrt{\log\left(\frac{2da}{\delta}\right)}/\tau_{n}\right)\geq 1-\frac{\delta}{2}\quad\forall\mathbf{x}\in\mathcal{X}_{n}.$ (39)
To cancel out the constants and keep the only dependence on stage $n$, we can set the discretization points $\tau_{n}=n^{2}brd\sqrt{\log\left(\frac{2da}{\delta}\right)}$ along each dimension of the feature space, leading to:
$P\left(\left|f(\mathbf{x})-f([\mathbf{x}]_{n})\right|\leq\frac{1}{n^{2}}\right)\geq 1-\frac{\delta}{2}\quad\forall\mathbf{x}\in\mathcal{X}_{n},$ (40)
where the total number of discretization points becomes $|\mathcal{X}_{n}|=\left(n^{2}brd\sqrt{\log\left(\frac{2da}{\delta}\right)}\right)^{d}$. Now, using $\frac{\delta}{2}$ in lemma 4.3 and choosing $\mathbf{x}=[\mathbf{x}]_{n}\in\mathcal{X}_{n}$ gives:
$\begin{split}\bigl|f([\mathbf{x}^{*}]_{n})-\mu_{n-1}([\mathbf{x}^{*}]_{n})\bigr|&=\bigl|f(\mathbf{x}^{*})-\mu_{n-1}([\mathbf{x}^{*}]_{n})+f([\mathbf{x}]_{n})-f(\mathbf{x}^{*})\bigr|\\[6.0pt] &\leq\bigl|f(\mathbf{x}^{*})-\mu_{n-1}([\mathbf{x}^{*}]_{n})\bigr|+\bigl|f([\mathbf{x}]_{n})-f(\mathbf{x}^{*})\bigr|\\[6.0pt] &\leq\tilde{\beta}_{n}^{1/2}\,\sigma_{n-1}([\mathbf{x}^{*}]_{n})+\frac{1}{n^{2}}.\end{split}$ (41)
The first inequality holds using triangle inequality, and the rest proceeds with probability $\geq 1-\delta$ after applying the union bound. Correspondingly, we have
$\tilde{\beta_{n}}=2\log\frac{|\mathcal{X}_{n}|\pi_{n}}{\delta/2}=2\log\frac{2\pi_{n}}{\delta}+2d\log\left(n^{2}brd\sqrt{\log\left(\frac{2da}{\delta}\right)}\right),$ (42)
which completes the proof. ∎
Lemma A.5.
For any arbitrarily small $\delta\in(0,1)$, choose $\tilde{\beta_{n}}=2\log\frac{4\pi_{n}}{\delta}+2d\log\left(n^{2}brd\sqrt{\log\left(\frac{4da}{\delta}\right)}\right),i.e.,$ $\beta_{n}\;=\;\frac{\sigma_{\min}(n)\,\Bigl[2\log\frac{4\pi_{n}}{\delta}+2d\log\left(n^{2}brd\sqrt{\log\left(\frac{4da}{\delta}\right)}\right)\Bigr]\;-\;A_{n}}{\sigma_{\min}(n)\;+\;B_{n}}\,.$ where $\sum_{n\geq 1}\pi_{n}^{-1}=1$, $\pi_{n}>0$. As $n\to\infty$, we have the following regret bound with probability $\geq 1-\delta$:
$r_{n}\leq 2d\tilde{\beta_{n}^{\frac{1}{2}}}\sigma_{n-1}(\mathbf{x}_{n}).$
Proof.
We start by choosing $\frac{\delta}{2}$ in both Lemmas 4.2 and 4.4, which implies that both lemmas will be satisfied with probability $\geq 1-\delta$. Choosing $\frac{\delta}{2}$ also gives
$\tilde{\beta_{n}}=2\log\frac{4\pi_{n}}{\delta}+2d\log\left(n^{2}brd\sqrt{\log\frac{4da}{\delta}}\right).$ (43)
Intuitively, it is a sensible choice as it is greater than the value of $\tilde{\beta_{n}}$ used in Lemma 4.4. Since the stage-$n$ location $\mathbf{x}_{n}$ is selected as the maximizer of the UCB metric, by definition we have:
$\mu_{n-1}(\mathbf{x}_{n})+\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n})\geq\mu_{n-1}([\mathbf{x}^{*}]_{n})+\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}([\mathbf{x}^{*}]_{n}).$ (44)
Applying Lemma 4.4 gives:
$\mu_{n-1}([\mathbf{x}^{*}]_{n})+\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}([\mathbf{x}^{*}]_{n})+\frac{1}{n^{2}}\geq f(\mathbf{x}^{*}).$ (45)
Combining all, we have:
$\mu_{n-1}(\mathbf{x}_{n})+\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n})\geq(1-d)\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}([\mathbf{x}^{*}]_{n})+f(\mathbf{x}^{*})-\frac{1}{n^{2}}.$ (46)
Thus,
$\begin{split}r_{t}=f(\mathbf{x}^{*})-f(\mathbf{x}_{n})&\leq\mu_{n-1}(\mathbf{x}_{n})+\tilde{\beta_{n}}^{\frac{1}{2}}\,\sigma_{n-1}(\mathbf{x}_{n})+\frac{1}{n^{2}}+\tilde{\beta_{n}}^{\frac{1}{2}}\,\sigma_{n-1}([\mathbf{x}^{*}]_{n}),\\ &\leq(d+1)\,\tilde{\beta_{n}}^{\frac{1}{2}}\,\sigma_{n-1}(\mathbf{x}_{n})+(d-1)\,\tilde{\beta_{n}}^{\frac{1}{2}}\,\sigma_{n-1}([\mathbf{x}^{*}]_{n})+\frac{1}{n^{2}}.\end{split}$ (47)
Since for all $\mathbf{x}\in\mathcal{X}$, we have $\lim_{n\to\infty}\|\mathbf{x}-[\mathbf{x}]_{n}\|=0$, suggesting that $\mathbf{x}^{*}]_{n}$ approaches $\mathbf{x}^{*}$ as $n$ increases to infinity. Plugging in, we have:
$r_{n}\leq 2\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n}).$ (48)
Lemma A.6.
The mutual information gain for a total of $T_{N}$} stages can be expressed as follows:
$I(y_{1:{T_{N}}};f_{1:{T_{N}}})=\frac{1}{2}\sum_{n=1}^{T_{N}}\log\left(1+\sigma^{-2}\sigma_{n-1}^{2}(\mathbf{x}_{n})\right)$ (49)
Proof.
Recall that $I(y_{1:{T_{N}}};f_{1:{T_{N}}})=H(y_{1:{T_{N}}})-\frac{1}{2}\log|2\pi e\sigma^{2}\mathbf{I}|$. Using the chain rule of conditional entropy gives:
$\begin{split}H\bigl(y_{1:T_{N}}\bigr)&=H\bigl(y_{1:T_{N-1}}\bigr)+H\bigl(y_{1:T_{N}}\mid y_{1:T_{N-1}}\bigr)\\[6.0pt] &=H\bigl(y_{1:T_{N-1}}\bigr)+\frac{1}{2}\,\log\!\bigl(2\pi e\,\bigl(\sigma^{2}+\sigma_{T_{N-1}}^{2}(\mathbf{x}_{T_{N}})\bigr)\bigr)\,.\end{split}$ (50)
Thus,
$\begin{split}I\bigl(y_{1:T_{N}};\,f_{1:T_{N}}\bigr)&=H\bigl(y_{1:T_{N-1}}\bigr)+\frac{1}{2}\log\bigl(2\pi e\,[\sigma^{2}+\sigma_{T_{N-1}}^{2}(\mathbf{x}_{T_{N}})]\bigr)-\frac{1}{2}\log\bigl|2\pi e\,\sigma^{2}\mathbf{I}\bigr|\\[6.0pt] &=H\bigl(y_{1:T_{N-1}}\bigr)+\frac{1}{2}\log\!\Bigl(1+\sigma^{-2}\,\sigma_{T_{N-1}}^{2}(\mathbf{x}_{T_{N}})\Bigr)\end{split}$ (51)
Note that $\mathbf{x}_{1},\ldots,\mathbf{x}_{T_{N}}$ are deterministic given the outcome observations $y_{1:T_{N-1}}$, and the conditional variance term $\sigma_{T_{N-1}}^{2}(\mathbf{x}_{T_{N}})$ does not depend on the realization of $y_{1:T_{N-1}}$ due to the conditioning property of the GP. The result thus follows by induction. ∎
Now we provide proof for the main theorem on the regret bound. We use $\mathcal{O}^{*}$, a variant of the $\mathcal{O}$ notation to suppress the log factors.
Proof.
Based on 4.5, we have $r_{n}^{2}\leq 2\tilde{\beta_{n}}\sigma_{n-1}^{2}(\mathbf{x}_{n})$ with probability $\geq 1-\delta$ as $n\to\infty$. Since $\tilde{\beta_{n}}=2\log\frac{4\pi_{n}}{\delta}+2d\log\left(n^{2}brd\sqrt{\log\frac{4da}{\delta}}\right)$ and is nondecreasing in $n$, we can upper bound it by the final stage $\tilde{\beta_{N}}$:
$2\tilde{\beta_{n}}\sigma_{n-1}^{2}(\mathbf{x}_{n})\leq 2\tilde{\beta_{N}}\sigma^{2}\frac{\sigma_{n-1}^{2}(\mathbf{x}_{n})}{\log(1+\sigma^{-2})}=(1/4)\tilde{\beta_{N}}C_{1}\log(1+\sigma^{-2}).$ (52)
Using Cauchy-Schwarz inequality, we have:
$\begin{split}R_{T_{N}}^{2}&\leq T_{N}\sum_{n=1}^{T_{N}}r_{n}^{2}\;\leq\;T_{N}\sum_{n=1}^{T_{N}}\frac{1}{4}\,\beta_{T_{N}}\,\log\bigl(1+\sigma^{-2}\bigr)\\[6.0pt] &=C_{1}\,T_{N}\,\beta_{T_{N}}\,I\bigl(y_{1:T_{N}};\,f_{1:T_{N}}\bigr)\\[6.0pt] &\leq C_{1}\,T_{N}\,\beta_{T_{N}}\,\gamma_{T_{N}}\,.\end{split}$ (53)
where $\gamma_{T_{N}}=\max I(y_{1:{T_{N}}};f_{1:{T_{N}}})$ is the maximum information gain after $T_{N}$} steps of sampling. Thus,
$P\left(R_{T_{N}}\leq\sqrt{C_{1}{T_{N}}\beta_{{T_{N}}}\gamma_{T_{N}}}\right)\geq 1-\delta.$ (54)
Note that our main theorem’s form is quite similar to {Srinivas et al., 2010}, although our stage-wise constant $\beta_{n,d}$ is different and includes a distance term.
A.3 Theorem 3: Two-Phase Regret
Proof Sketch for the Two-Phase Regret Decomposition
Let $N$ be the total number of rounds. Suppose the first $T_{M}$} rounds use the full GP-UCB strategy, while rounds $t={T_{M+1}}$ to $t={T_{N}}$ employ a GSSBO strategy. The cumulative regret is denoted by $R_{T_{N}}\;=\;\sum_{t=1}^{T_{N}}\bigl(f(\mathbf{x}^{*})-f(\mathbf{x}_{t})\bigr)$ , where $\mathbf{x}^{*}$ is an optimal point and $\mathbf{x}_{t}$ is the decision made at time $t$. Decompose the $N$ rounds into two segments:
$R_{T_{N}}\;=\;\underbrace{\sum_{t=1}^{{T_{M}}}\bigl(f(\mathbf{x}^{*})-f(\mathbf{x}_{t})\bigr)}_{\displaystyle R_{{T_{M}}}^{(\text{full})}}\;+\;\underbrace{\sum_{t={T_{M}}+1}^{{T_{N}}}\bigl(f(\mathbf{x}^{*})-f(\mathbf{x}_{t})\bigr)}_{\displaystyle R_{{T_{N}}-{T_{M}}}^{(\text{selected})}}.$ (55)
1. Regret Bound in the First $T_{M}$} Rounds
During the initial $M$ rounds, the strategy relies on the standard GP-UCB. By the well-known GP-UCB regret bounds [Srinivas et al., 2009], there exists a constant $C_{1}=\frac{8}{\log(1+\sigma^{-2})}$, pick $\delta\in(0,1)$, and define $\beta_{n}=2\log\left(n^{2}2\pi^{2}/(3\delta)\right)+2d\log\left(n^{2}brd\sqrt{\log(4da/\delta)}\right)$, we have,
$\Pr\left\{R_{M}^{(\text{full})}\leq\sqrt{C_{1}M\beta_{M}\gamma_{M}}+2\quad\forall M\geq 1\right\}\geq 1-\delta.$ (56)
2. Regret Bound from Round ${T_{M}}+1$ to $T_{N}$}
Starting from iteration $t={T_{M}}+1$, the regret analysis switches to a sparse GP-UCB. Let $R_{{T_{N}}-{T_{M}}}^{(\text{selected})}$ denote the regret incurred in these final ${T_{N}}-{T_{M}}$ rounds. Choose $\beta_{n}\;=\;\frac{\sigma_{\min}(n)\,\Bigl[2\log\frac{4\pi_{n}}{\delta}+2d\log\left(n^{2}brd\sqrt{\log\left(\frac{4da}{\delta}\right)}\right)\Bigr]\;-\;A_{n}}{\sigma_{\min}(n)\;+\;B_{n}}\,.$ where $\sum_{n\geq 1}\pi_{n}^{-1}=1,\quad\pi_{n}>0$. As $n\to\infty$, we obtain a regret bound of $\mathcal{O}^{*}(\sqrt{d({T_{N}}-{T_{M}})\gamma_{({T_{N}}-{T_{M}})}})$. Specifically, we have:
$\Pr\left(R_{{T_{N}}-{T_{M}}}\leq\sqrt{C_{1}({T_{N}}-{T_{M}})\beta_{({T_{N}}-{T_{M}})}\gamma_{({T_{N}}-{T_{M}})}}\quad\forall{({T_{N}}-{T_{M}})}\geq 1\right)\geq 1-\delta.$ (57)
3. Overall Regret
Summarizing both phases, the total regret satisfies $R_{T_{N}}\;=\;R_{{T_{M}}}^{(\text{full})}\;+\;R_{{T_{N}}-{T_{M}}}^{(\text{selected})}$
Consequently,
$\Pr\left(R_{{T_{N}}}\leq\sqrt{\,C_{1}\,{T_{M}}\,\beta_{T_{M}}\,\gamma_{T_{M}}}+2+\sqrt{\,C_{1}\,({T_{N}}-{T_{M}})\,\beta_{({T_{N}}-{T_{M}})}\,\gamma_{({T_{N}}-{T_{M}})}}\quad\forall{{T_{N}}}\geq 1\right)\geq 1-\delta.$ (58)
Appendix B Appendix Experiments
B.1 Supplementary experiments
Figure 5 and 6 compares the cumulative runtime over $1000$ iterations on Eggholder2, Levy20, Rastrigin100 functions and NAS experiment (in seconds).
Figure 5: Cumulative time cost of algorithms 2.
Figure 6: Cumulative time cost of algorithms 3.
B.2 Subset Samples Distribution Study
Figure 7 illustrates the sample distribution of GSSBO and Standard GP-UCB, on the first two dimensions of the Hartmann6 function. In this experiment, we recorded the first 200 sequential samples from a standard BO process and constrained the buffer size to 100. The objective is to identify the global minimum, and darker-colored samples correspond to values closer to the optimal. During the optimization process, only a small number of samples are located near the optimal value. As shown in the middle panel, the gradient-based sample selection method selects a more informative and diverse subset, retaining more samples closer to the optimal or suboptimal, which is indicated by preserving a higher number of darker-colored samples in the figure. In contrast, the random selection strategy reduces the sample density uniformly across all regions, leading to a significant loss of samples near the optimal or suboptimal, as represented by the retention of many lighter-colored samples in the right panel.
While BO algorithms are theoretically designed to balance exploitation and exploration, with limited budget in practice, they can over-exploit current best regions before shifting to exploration [Wang and Ng, 2020], leading to suboptimal performance in locating the global optimums. With gradient-based sample selection, the relative density of samples near the optimal and suboptimal regions increases, maintaining a more balanced distribution. This subset encourages subsequent iterations to focus on regions outside the optimal and suboptimal regions, promoting the exploration of other regions of the search space. As a result, the over-exploitation issue is mitigated.
Figure 7: Sample distribution: GP-UCB (left), GSSBO (middle), and RSSBO (right).
B.3 Experiments on Kmeans++ selection
Oglic [Oglic and Gärtner, 2017] and Hayakawa [Hayakawa et al., 2023] proposed to select a subset in RKHS, then employed them to construct the Nyström low-rank approximation. We included it as an additional baseline, experimental results in Fig.8 demonstrate that K-means++ has no clear advantage in cumulative regret compared to our proposed method in subset selection, and it costs more time than our method.
Figure 8: Cumulative regret of GSSBO and Kmeans++ on the Eggholder2, Hart6, Levy20, Powell50, Rastrigin100 functions and NAS experiment.
B.4 Experiments on high-dimensional BO methods
We noticed that all these algorithms struggled with high-dimensional tasks, so we included some high-dimensional baselines: REMBO [Wang et al., 2016] and HESBO [Nayebi et al., 2019]. The results are shown in Fig. 9, which shows that the difference between the three methods is subtle on Levy20, but REMBO and HSEBO do not perform well on higher-dimensional tasks because these two dimensionality reduction methods do not use all the dimensional information.
Figure 9: Cumulative regret of GSSBO, REMBO, and HSEBO on the Levy20, Powell50, Rastrigin100 functions.
B.5 Experiments on other surrogates
Although our theoretical analysis focuses on Gaussian Process surrogates, our proposed gradient-based sample selection method is not limited to GPs. The core idea of the gradient-based sample selection method is to use gradient information to measure the importance of each sample to model fitting, which is model-independent. Therefore, this method can also be applied to Bayesian neural networks (BNN) [Mullachery et al., 2018] and Deep Kernel(DK) methods [Wilson et al., 2016]. We have extended our experiments to other popular surrogate models such as Deep Kernels and Bayesian Neural Networks. Our additional results demonstrate that the proposed sample selection strategy yields performance improvements across these different surrogate models. The updated experimental results can be found in Fig. 10.
Figure 10: Cumulative regret of Deep Kernal and BNN with UCB algorithms on the Hart6 experiment.
B.6 RMSE Analysis of Subset GP Fits
To validate that the selected subset is indeed “representative and informative”, we compare the predictive quality of three surrogate models in terms of RMSE on a test dataset: (i) the full-data GP used in GP-UCB, (ii) the GP trained on the gradient-based subset (GSSBO), and (iii) the GP trained on a randomly chosen subset of the same size (RSSBO). Starting from the 60th iteration, the subset selection algorithm is used. At each iteration, we construct the corresponding GP surrogate and compute the root mean square error (RMSE) between its posterior mean and the true underlying function values on the test set:
$\text{RMSE}=\sqrt{\frac{1}{T}\sum_{i=1}^{T}\left(\mu(x_{i})-f(x_{i})\right)^{2}}.$ (59)
As shown in the figure11, GSSBO consistently achieves RMSE values that are close to, and in many iterations only marginally worse than, the full-data GP, indicating that the gradient-selected subset preserves most of the essential information for accurate function approximation. In contrast, RSSBO exhibits notably higher error and larger variance over time, reflecting that random subset sampling fails to reliably capture the non-redundant information necessary for stable surrogate quality. The fluctuations in all curves are partially due to observation noise and the dynamic subset update process, but GSSBO maintains a relatively low and stable RMSE throughout, which empirically supports our design of combining gradient-based importance with direction-space diversity to build a compact yet informative subset that effectively approximates the full GP.
Figure 11: Cumulative regret of Deep Kernal and BNN with UCB algorithms on the Hart6 experiment.